OpenStudy (anonymous):

A manufacturer finds that the revenue generated by selling x gadgets is given by the function R(x)=320x-.2x^2, where the revenue r(x) is measurd in dollars. What is the maximum revenue?

OpenStudy (anonymous):

can anyone do this please

OpenStudy (anonymous):

Just take the derivative, then solve for when the derivative equals zero. Then plug that solution (s) into your original equation -.2(2)x+320 = 0 x=320/ 2(.2) max= 320(320/(2(.2)) - .2 ( 320/ (2(.2)))^2